OpenStudy (anonymous):

I WILL ABSOLUETLY GIVE YOU A MEDAL AND FAN IF YOU HELP ME GET THIS ANSWER, ALMOST DONE PLEASE!!! The diameter of a cylinder is 8 cm and its height is 14 cm. Calculate the volume of the cylinder to the nearest cubic centimetre. Use π = 3.14.

OpenStudy (anonymous):

Thankyou! Please help!

OpenStudy (thehollywoodstud):

3562.57

Nnesha (nnesha):

what's the formula to find volume of cylinder ??

OpenStudy (thehollywoodstud):

V=πr2h

OpenStudy (anonymous):

IDK!

OpenStudy (thehollywoodstud):

volume =pi time radius squared times height

OpenStudy (anonymous):

v=pie*(r^2)*h v= 3.14*4^2*18 v=3.14*16*18 v=904.32cubic unit

OpenStudy (thehollywoodstud):

v=3.14r^2h

Nnesha (nnesha):

\[\huge\rm v = \pi r^2 h\] remember radius is half of diameter \[\large\rm r = \frac{ d }{ 2 }\] replace d by 8 and then replace r and h by their values

OpenStudy (anonymous):

the actual question is one I have to write in, so I have no options, so I just need the accurate answer

OpenStudy (thehollywoodstud):

703.72

OpenStudy (anonymous):

this is for sure the answer??

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @EmilyBunny45 the actual question is one I have to write in, so I have no options, so I just need the accurate answer \(\color{blue}{\text{End of Quote}}\) no need for A.C replace variables by their values

OpenStudy (anonymous):

Check my answer

OpenStudy (thehollywoodstud):

yes the diameter is 8 which means the radius is 4

OpenStudy (anonymous):

904.32?

OpenStudy (anonymous):

yh.

OpenStudy (thehollywoodstud):

no 703.72

OpenStudy (anonymous):

I am confused now

OpenStudy (thehollywoodstud):

4^2=16*14is 224 and then that time pi is 703.72

OpenStudy (anonymous):

(3.14* 4^2)14=703.36

OpenStudy (thehollywoodstud):

i agree with robtobey

OpenStudy (thehollywoodstud):

sorry i meant 703.36

OpenStudy (anonymous):

Sorry the and is 703.36. I accidentally used 18 as height. Thehollywoodstud is right

OpenStudy (anonymous):

okay thanks everyone, I can only give one medal, but I will fan y'all

OpenStudy (thehollywoodstud):

703.36

OpenStudy (anonymous):

For method you can refer mine, but change the height in the one I used.

OpenStudy (thehollywoodstud):

if u need more help you can ask me